A stationary store has decided to accept a large shipment of ball-point pens if an inspection of 19 randomly selected pens yields no more than two defective (a) Find the probability that this shipment is accepted if 1(7% of the total shipment is defective (use 3 decirnal places) 0.706 b) Find the probabity that this shipment is not accepted if 20% of the total shipment is defective. (use 3 decimal places.) 0.764
Answer #1

Solution:-

b) The probability that this shipment is not accepted if 20% of the total shipment is defective is 0.763.

n = 19, p = 0.20

x = 2

By applying binomial distribution:-

P(x,n) = nCx*px*(1-p)(n-x)

P(x > 2) = 1 - [P(x = 0) + P(x = 1) + P(x = 2)]

P(x > 2) = 0.7631